Identify what your variables represent. Write two equations and solve the system.
After a snow advisory in January, Fleet Farm had a run on snow shovels. They had the Snowpusher on sale for $15.95 and the lighter weight Snowdozer on sale for $7.95. By the end of one week, t he entire inventory of 415 shovels had been sold out and the receipts for them totaled $4779.25. How many of each type did they sell?

Answer by jorel1380(3719)

x+y=415
15.95x+7.95y=4779.25
------------------
1595x+795y=477925
795x+795y=329925
800x=148000
x=185
y=230
185 Snowpushers and 230 Snowdozers were sold for the week..